What is the mass (in grams) of 10.00 L of propane vapor (C₃H₈) at STP?


1
Expert's answer
2021-04-07T11:32:29-0400

At STP condition: 1 mole of a gas= 22.4 L

moles of C3H8 = 10.0L/22.4L = 0.446 mol

Mass of C3H8 = 0.446×44.1 = 19.668 g